OpenStudy (anonymous):

Simplify. 4 square root of 2 end root plus 7 square root of 2 end root minus 3 square root of 2

OpenStudy (anonymous):

\[4\sqrt{2}+7\sqrt{2}-3\sqrt{2}\]